how would you Determine whether each triangle is a right triangle.
28. c=34,a=16,b=30


Answer by checkley77(12844)

a^2+b^2=c^2 where c is the longest (hypotenuse) side.
16^2+30^2=34^2
256+900=1,156
1,156=1,156 this is a right triangle.